OpenStudy (anonymous):

Solve the system y = 3x + 2 and 3y = 9x + 6 would this be no solution?

OpenStudy (anonymous):

yes it is no solution because they are both equal to each other... the lines pretty much overlay each other.

OpenStudy (anonymous):

they r equal!!

OpenStudy (anonymous):

plug first equation in second equation 3(3x+2)=9x+6 9x+6=9x+6 so, these two equation represent same equation therefore we have infinite number of solution

OpenStudy (anonymous):

answer is infinite number of solution

OpenStudy (anonymous):

let , x = k ( any variable ) so, y = 3k+2 where k be any real number

OpenStudy (anonymous):

so,we can get infinite number,as x1/x2=y1/y2=c1/c2

OpenStudy (anonymous):

they are the same